Tatu Corp. makes all sales on credit. During May 2016, total credit sales were $2,650,000, collections were $2,400,000 and ac with solution please
Answer #1
Amount Percent Allowance required
a b a*b
Due within 10 days 20000 1% 200
Past due 30-90 days 10000 10% 1000
Past due more than 90 days 10000 30% 3000
Total 4200
Allowance for baddebt as on May 31,2016=Allowance for baddebt as on May 1,2016-Accounts written off as uncollectible=3000-25000=-22000=$ 22000 debit
Required balance in allowance for baddebt=$ 4200 credit
Baddebt expense=Allowance for baddebt as on May 31,2016+Required balance in allowance for baddebt=22000+4200=$ 26200
Answer is
B. $26200